# Approximation in Hölder spaces

### 03 Jan 2024 v1

@@ 1.2. Applications on the Euclidean space
$\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n};Y)=\overline{C^{\infty}_{c}(\mathbb{R}^{n};Y)}^{\dot{C}^{0,\omega}(\mathbb{R}^{n};Y)}.$
To make a connection with the mean oscillation conditions, we next set some definitions. For the sake of simplicity, we now consider $X=\mathbb{R}^{n}$ to be the Euclidean space, but $Y$ is still allowed to be an arbitrary Banach space. We first define the bounded and vanishing mean oscillation spaces with a general continuity moduli.
@@
Let $\omega$ be a modulus and $Y$ a Banach space and define
@@
One of the most interesting results in Guo et al. is their [GHWY21, Theorem 1.7.] stating that $\mathrm{VMO}^{\alpha}(\mathbb{R}^{n})=\overline{C^{\infty}_{c}(\mathbb{R}^{n})}^{\mathrm{BMO}^{\alpha}(\mathbb{R}^{n})};$ the inclusion ``""```$\supset$"` ` ⊃ " is immediate, while ``"``"`$\subset$"` ` ⊂ " is completely not so. (We also note that notationally $\operatorname{CMO}=\mathrm{VMO}$ for Guo et al.) Also, if $\alpha\geq 1,$ then $\mathrm{VMO}^{\alpha}(\mathbb{R}^{n})$ consists of the constant functions, simply by noting that $\mathrm{VMO}^{\alpha}_{\operatorname{small}}(\mathbb{R}^{n})$ forces the Lipschitz constant to be zero. Thus, only the range $\alpha\in[0,1)$ needs to be considered. Importantly, for $\alpha=0,$ the approximation $\mathrm{VMO}(\mathbb{R}^{n})=\overline{C^{\infty}_{c}(\mathbb{R}^{n})}^{\mathrm{BMO}(\mathbb{R}^{n})}$ was already present in the fundamental work of Uchiyama [Uch1978]. The approximation of Uchiyama does not work when $\alpha>0$ and Guo et al. [GHWY21] constructed a non-trivial new approximation, which depends crucially on the existence of a dyadic grid on the underlying space $\mathbb{R}^{n}$ and a careful construction on top of this grid. One fact of our result 1.10, and in particular of its proof, is that it provides a conceptually easier approximation using only basic mollification and a (somewhat) basic truncation.
To make the connection with the mean oscillations of [GHWY21, Theorem 1.7.] and the pointwise conditions of our Theorem 1.10 we impose the following
$\|\omega\|_{\operatorname{Dini}}:=\sup_{s>0}\frac{1}{\omega(s)}\int_{0}^{s}\omega(t)\frac{\,\mathrm{d}t}{t}<\infty$ (6)
multiscale Dini-type summability condition.
@@
$\frac{\|\cdot\|_{\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}}{\|\cdot\|_{\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)}}+\frac{\|\cdot\|_{\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)}}{\|\cdot\|_{\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}}\lesssim_{n,C_{\operatorname{db}}}\|\omega\|_{\operatorname{Dini}}.$ (7)
Moreover, for each $\Upsilon\in\{\operatorname{small},\operatorname{far},\operatorname{large}\},$ there holds that
$\dot{\operatorname{VC}}^{0,\omega}_{\Upsilon}(\mathbb{R}^{n},Y)=\mathrm{VMO}^{\omega}_{\Upsilon}(\mathbb{R}^{n},Y),\qquad\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y)=\mathrm{VMO}^{\omega}(\mathbb{R}^{n},Y).$ (8)
The proof-idea of Theorem 1.13 of linking a pointwise difference with a telescoping (martingale) sum of differences of averages traces back to Meyers [Mey1964], who proved a version of (7) with the modulus $t^{\alpha},$ for $\alpha>0.$ To our best knowledge, the condition (6) and the identification (8) are both new.
Now combining our Theorems 1.10 and 1.13, we obtain the following Theorem 1.14 generalizing the approximation [GHWY21, Theorem 1.7.] of Guo et al.
@@
$\mathrm{VMO}^{\omega}(\mathbb{R}^{n};Y)=\overline{C^{\infty}_{c}(\mathbb{R}^{n};Y)}^{\mathrm{BMO}^{\omega}(\mathbb{R}^{n};Y)}.$
@@ 2. Bounded support approximation
In this section prove our first main Theorem 1.5 and then we show how to easily obtain from this Theorem 1.10. At the end of this section, we also prove Theorem 1.13. Recall that we simplify notation by denoting $\|x\|_{X}=\|x\|$ and $\|f(x)\|_{Y}=|f(x)|.$ We begin with two simple lemmas.
@@
Let $f\in\dot{\operatorname{VC}}^{0,\omega}$ and $\varepsilon>0.$ By Lemma 2.1 we find two parameters $0<r\ll R$ such that
@@
Now, let us denote $g:=f\circ\tau\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}$ and let $r>0$ be such that
$\|g\|_{\dot{C}^{0,\omega}_{r}}:=\sup_{\begin{subarray}{c}x\not=y\in X\\ \|x-y\|<r\end{subarray}}\frac{|g(x)-g(y)|}{\omega(\|x-y\|)}\leq\varepsilon.$ (26)
By (15), the fact that $\operatorname{Lip}(\tau)\leq 5$ (see (19)), and the proof of Lemma 2.2, this new $r>0$ can be taken to be an absolute multiple of the one in (15). Therefore, we may and do assume that both (15) and (26) hold for the same $r=r(\varepsilon),$ which is of course independent of $R$ and $M.$ Next we show that
@@
Provided that $M=M(\varepsilon,R)$ is large enough, the condition $\omega(\infty)=\infty$ says that the last term can be made smaller than $\varepsilon\omega(M^{1/4})\leq\varepsilon\omega(\|x-z\|),$ and thus (27) holds for these $x$ and $z.$ Therefore, during the rest of the verification of (27) we can assume that $\|x-z\|\leq M^{1/4},$ whenever needed. Notice also that in the case $x,z\in X\setminus B(0,2M)$ there is nothing to prove, as then $g(x)=g(z)=0.$ Thus, we can assume that botj $x,z\in B(0,10M).$ In this case we bound
@@
$\sum_{k=0}^{\infty}\omega(\ell(Q_{k}(x)))=\sum_{k=0}^{\infty}\ell(Q_{k}(x))\frac{\omega(\ell(Q_{k}(x)))}{\ell(Q_{k}(x))}\leq 2\int_{0}^{\ell(Q_{0})}\frac{\omega(t)}{t}\,\mathrm{d}t\lesssim\|\omega\|_{\operatorname{Dini}}\omega(\ell(Q_{0})),$
and, since $\omega(\ell(Q_{0}))=\omega(|x-y|),$ we have shown that $\|f\|_{\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}\lesssim_{n}\|\omega\|_{\operatorname{Dini}}\|f\|_{\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)}.$
It is clear that the above argument gives $\dot{\operatorname{VC}}^{0,\omega}_{\Upsilon}(\mathbb{R}^{n},Y)=\mathrm{VMO}^{\omega}_{\Upsilon}(\mathbb{R}^{n},Y)$ when $\Upsilon=\operatorname{small},$ as well as the inclusion $\dot{\operatorname{VC}}^{0,\omega}_{\Upsilon}(\mathbb{R}^{n},Y)\subset\mathrm{VMO}^{\omega}_{\Upsilon}(\mathbb{R}^{n},Y)$ for $\Upsilon=\operatorname{far}.$ The reverse inclusion ``""```$\supset$"` ` ⊃ " for $\Upsilon=\operatorname{far}$ can be seen as follows. Let $R>0$ be such that if $\operatorname{\mathit{d}}(Q,0)>R,$ then $\mathcal{O}^{\omega}(f,Q)\leq\varepsilon.$ Let $|x|,|y|>R$ both be far from the origin. Using the structure of the Euclidean space, clearly there exists a dimensional constant $N(n)$ and points and cubes $\{z_{k},Q_{k}\}_{k=1}^{N}$ such that $z_{0}=x,$ $z_{N}=y$ and $z_{k},z_{k+1}\in Q_{k}$ and $\operatorname{\mathit{d}}(Q_{k},0)\geq R$ and $\ell(Q_{k})=|z_{k}-z_{k+1}|.$ Thus, by the chain of inequalities that led us to the right estimate for (32), we obtain
$|f(x)-f(y)|_{Y}\leq\sum_{k=1}^{N}|f(z_{k})-f(z_{k+1})|_{Y}\lesssim\|\omega\|_{\operatorname{Dini}}\varepsilon\sum_{k=1}^{N}\omega(\ell(Q_{k}))\lesssim\varepsilon\omega(|x-y|),$
which concludes the proof of the inclusion $\dot{\operatorname{VC}}^{0,\omega}_{\Upsilon}(\mathbb{R}^{n},Y)\supset\mathrm{VMO}^{\omega}_{\Upsilon}(\mathbb{R}^{n},Y).$
For $\Upsilon=\operatorname{large},$ a further argument is needed for both inclusions. We begin by showing that $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y)\supset\mathrm{VMO}^{\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y).$ Let $f\in\mathrm{VMO}^{\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y),$ and $\varepsilon>0,$ and let $K\in\mathbb{N}$ be such that if $\ell(Q)\geq 2^{K},$ then $\mathcal{O}^{\omega}(f;Q)\leq\varepsilon.$ Let $N\gg K$ be a large integer to be specified later, and suppose $x,y\in\mathbb{R}^{n}$ are so that $|x-y|\geq 2^{N}.$ Let $M\geq N$ be so that $2^{M}\leq\left|x-y\right|\leq 2^{M+1},$ and let $Q_{0}$ be a cube containing both $x,y$ and with $\ell(Q_{0})=|x-y|.$ Then, again considering the left sum in the expansion (32), we bound
@@
$\lesssim\Big{[}\varepsilon\|\omega\|_{\operatorname{Dini}}+\frac{\omega(2^{K})}{\omega(2^{N+1})}\|f\|_{\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)}\|\omega\|_{\operatorname{Dini}}\Big{]}\omega(|x-y|).$
By the condition $\omega(\infty)=\infty$, we choose $N$ large enough and bound the bracketed term from above by $\leq 2\varepsilon\|\omega\|_{\operatorname{Dini}}.$ Repeating the same proof with $y$ in place of $x$, we control the right sum on the line (32) and thus obtain $|f(x)-f(y)|\leq 4\varepsilon\|\omega\|_{\operatorname{Dini}}\omega(|x-y|),$ that is, $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y).$ Then, we show that $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y)\subset\mathrm{VMO}^{\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y).$ Let $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y)$, $\varepsilon>0,$ and $R>0$ be such that $\operatorname{osc}^{\omega}_{\delta}(f)\leq\varepsilon,$ if $\delta\geq R.$ Let $M\gg R$ be a large constant, which we will specify later, and $Q\in\mathcal{Q}$ a cube such that $\ell(Q)\geq M.$ For every $y\in Q,$ denote by $Q_{R}(y)$ the cube centered at $y$ and of $\operatorname{diam}(Q_{R}(y))=R.$ Then, we have
@@ 3. Approximation in infinite dimensional spaces
In this section we study smooth approximations in Banach spaces with respect to the $\dot{C}^{0,\omega}$ seminorm. In Subsections 3.1 and 3.2 we establish results for real-valued functions $f:X\to\mathbb{R}$, where the approximating functions are not only $C^{k}$ or $C^{1,\alpha}$ smooth but also Lipschitz and with bounded support. In the last Subsection 3.3 we obtain $C^{\infty}$ smooth approximations of Banach-valued mappings $f:c_{0}(\Gamma)\to Y,$ that are not necessarily Lipschitz, but have bounded support and belong to $\dot{\operatorname{VC}}^{0,\omega}(X,Y).$ As a corollary, we deduce $C^{\infty}$ and $\dot{\operatorname{VC}}^{0,\omega}$ and approximations with bounded support for all $\dot{\operatorname{VC}}^{0,\omega}$ functions defined on $\mathbb{R}^{n}$ and with values in any Banach space.

### 21 Feb 2024 v2

@@ 1.2. Applications on $X=\mathbb{R}^{n}$
$\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y)=\overline{C^{\infty}_{c}(\mathbb{R}^{n},Y)}^{\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}.$
To connect our pointwise conditions with vanishing mean oscillation type conditions and to establish the two descriptions of these spaces (one through approximations and one through vanishing mean oscillations) we next set some definitions. For simplicity, we let $X=\mathbb{R}^{n}$ be the Euclidean space, but $Y$ is still allowed to be an arbitrary Banach space.
@@
Let $\omega$ be a modulus and $Y$ a Banach space and define the vanishing scales
@@
One of the most interesting results in Guo et al. [GHWY21, Theorem 1.7.] states that $\mathrm{VMO}^{\alpha}(\mathbb{R}^{n})=\overline{C^{\infty}_{c}(\mathbb{R}^{n})}^{\mathrm{BMO}^{\alpha}(\mathbb{R}^{n})},$ for $\alpha\in(0,1);$ this is exactly the equivalence between approximations and vanishing mean oscillations mentioned above. (We also note that $\operatorname{CMO}^{\alpha}=\mathrm{VMO}^{\alpha}$ for Guo et al.) Importantly for $\alpha=0,$ the approximation $\mathrm{VMO}(\mathbb{R}^{n})=\overline{C^{\infty}_{c}(\mathbb{R}^{n})}^{\mathrm{BMO}(\mathbb{R}^{n})}$ was already present in the work of Uchiyama [Uch1978]. Uchiyama’s construction does not easily translate to the case $\alpha\in(0,1),$ and Guo et al. [GHWY21] provided a non-trivial new approximation that depends on the existence of a dyadic grid on the underlying space $\mathbb{R}^{n}$ (as does Uchiyama’s original approximation) and a careful construction using this grid. One fact of our result 1.8, and in particular of its proof, is that it provides a conceptually easier approximation using only basic mollification and a (somewhat) basic truncation of the support.
Finally, to make the connection with the mean oscillations of [GHWY21, Theorem 1.7] and the pointwise conditions of Theorem 1.8 we assume the following summability condition
$[\omega]_{*}:=\sup_{s>0}\frac{1}{\omega(s)}\int_{0}^{s}\omega(t)\frac{\,\mathrm{d}t}{t}<\infty.$ (7)
@@
$*.\displaystyle\frac{\|\cdot\|_{\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}}{\|\cdot\|_{\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)}}+\frac{\|\cdot\|_{\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)}}{\|\cdot\|_{\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}}\lesssim_{n,C_{\operatorname{db}}}[\omega]_{*}.$ (8)
Moreover, for each $\Gamma\in\{\operatorname{small},\operatorname{far},\operatorname{large}\},$ there holds that
$\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}(\mathbb{R}^{n},Y)=\mathrm{VMO}^{\omega}_{\Gamma}(\mathbb{R}^{n},Y),\qquad\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y)=\mathrm{VMO}^{\omega}(\mathbb{R}^{n},Y).$ (9)
The proof-idea of Theorem 1.11 of linking a pointwise difference with a telescoping sum of differences of averages traces back to Meyers [Mey1964], who proved a version of (8) with the modulus $t^{\alpha},$ for $\alpha>0.$ It seems to us that the condition (7) is new, but it is natural summability condition, after all. To our best knowledge, the identification (9) is new. Combining Theorems 1.8 and 1.11, we obtain the following.
@@
$\mathrm{VMO}^{\omega}(\mathbb{R}^{n},Y)=\overline{C^{\infty}_{c}(\mathbb{R}^{n},Y)}^{\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)}.$
@@ 2. Bounded support approximation
In this section we prove Theorem 1.4 and then show how to easily obtain from this Theorem 1.8. At the end of this section, we also prove Theorem 1.11. Recall that we simplify notation by denoting $\|x\|_{X}=\|x\|$ and $\|f(x)\|_{Y}=|f(x)|.$ We begin with two simple lemmas.
@@
First, note that if $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)$ and has bounded support, then, by property (3) of $\omega,$ $f$ belongs to $\dot{\operatorname{VC}}^{0,\omega}(X,Y).$ Since $\dot{\operatorname{VC}}^{0,\omega}(X,Y)$ is closed under limits with respect to the seminorm $\dot{C}^{0,\omega},$ the inclusion $\overline{\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)\cap C_{\operatorname{bs}}(X,Y)}\subset\dot{\operatorname{VC}}^{0,\omega}(X,Y)$ holds true.
To prove the reverse (much more complicated) inclusion, let $f\in\dot{\operatorname{VC}}^{0,\omega}$ and $\varepsilon>0.$ By Lemma 2.1 we find two parameters $0<r\ll R$ such that
@@
Now, let us denote $g:=f\circ\tau\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}$ and let $r^{\prime}>0$ be such that
$\|g\|_{\dot{C}^{0,\omega}_{r^{\prime}}}:=\sup_{\begin{subarray}{c}x\not=y\in X\\ \|x-y\|<r^{\prime}\end{subarray}}\frac{|g(x)-g(y)|}{\omega(\|x-y\|)}\leq\varepsilon.$ (27)
By (16), the fact that $\operatorname{Lip}(\tau)\leq 5$ (see (20)), and the proof of Lemma 2.2, this new $r^{\prime}>0$ can be taken to be an absolute multiple of the $r$ in (16). Therefore, we may and do assume that both (16) and (27) hold for the same $r=r(\varepsilon),$ which is of course independent of $R$ and $M.$
Next we show that
@@
Provided that $M=M(\varepsilon,R)$ is large enough, the condition $\omega(\infty)=\infty$ says that the last term can be made smaller than $\varepsilon\omega(M^{1/4})\leq\varepsilon\omega(\|x-z\|),$ and thus (28) holds for these $x$ and $z.$ Therefore, during the rest of the verification of (28) we can assume that $\|x-z\|\leq M^{1/4},$ whenever needed. Notice also that in the case $x,z\in X\setminus B(0,2M)$ there is nothing to prove, as then $g(x)=g(z)=0.$ Thus, we can assume that both $x,z\in B(0,10M).$ In this case we bound
@@
$\sum_{k=0}^{\infty}\omega(\ell(Q_{k}(x)))=\sum_{k=0}^{\infty}\ell(Q_{k}(x))\frac{\omega(\ell(Q_{k}(x)))}{\ell(Q_{k}(x))}\leq 2\int_{0}^{\ell(Q_{0})}\frac{\omega(t)}{t}\,\mathrm{d}t\lesssim[\omega]_{*}\omega(\ell(Q_{0})),$
and, since $\omega(\ell(Q_{0}))=\omega(|x-y|),$ we have shown that $\|f\|_{\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}\lesssim_{n}[\omega]_{*}\|f\|_{\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)}.$
It is clear that the above argument gives $\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}(\mathbb{R}^{n},Y)=\mathrm{VMO}^{\omega}_{\Gamma}(\mathbb{R}^{n},Y)$ when $\Gamma=\operatorname{small},$ as well as the inclusion $\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}(\mathbb{R}^{n},Y)\subset\mathrm{VMO}^{\omega}_{\Gamma}(\mathbb{R}^{n},Y)$ for $\Gamma=\operatorname{far}.$ The reverse inclusion ``""```$\supset$"` ` ⊃ " for $\Gamma=\operatorname{far}$ can be seen as follows. Let $R>0$ be such that if $\operatorname{\mathit{d}}(Q,0)>R,$ then $\mathcal{O}^{\omega}(f,Q)\leq\varepsilon.$ Let $|x|,|y|>R$ both be far from the origin. Using the structure of the Euclidean space, clearly there exists a dimensional constant $N(n)$ and points and cubes $\{z_{k},Q_{k}\}_{k=1}^{N}$ such that $z_{0}=x,$ $z_{N}=y$ and $z_{k},z_{k+1}\in Q_{k}$ and $\operatorname{\mathit{d}}(Q_{k},0)\geq R$ and $\ell(Q_{k})=|z_{k}-z_{k+1}|.$ Thus, by the chain of inequalities that led us to the right estimate for (33), we obtain
$|f(x)-f(y)|_{Y}\leq\sum_{k=1}^{N}|f(z_{k})-f(z_{k+1})|_{Y}\lesssim[\omega]_{*}\varepsilon\sum_{k=1}^{N}\omega(\ell(Q_{k}))\lesssim\varepsilon\omega(|x-y|),$
which concludes the proof of the inclusion $\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}(\mathbb{R}^{n},Y)\supset\mathrm{VMO}^{\omega}_{\Gamma}(\mathbb{R}^{n},Y).$
For $\Gamma=\operatorname{large},$ a further argument is needed for both inclusions. We begin by showing that $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y)\supset\mathrm{VMO}^{\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y).$ Let $f\in\mathrm{VMO}^{\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y),$ and $\varepsilon>0,$ and let $K\in\mathbb{N}$ be such that if $\ell(Q)\geq 2^{K},$ then $\mathcal{O}^{\omega}(f;Q)\leq\varepsilon.$ Let $N\gg K$ be a large integer to be specified later, and suppose $x,y\in\mathbb{R}^{n}$ are so that $|x-y|\geq 2^{N}.$ Let $M\geq N$ be so that $2^{M}\leq\left|x-y\right|\leq 2^{M+1},$ and let $Q_{0}$ be a cube containing both $x,y$ and with $\ell(Q_{0})=|x-y|.$ Then, again considering the left sum in the expansion (33), we bound
@@
$\lesssim\Big{[}\varepsilon[\omega]_{*}+\frac{\omega(2^{K})}{\omega(2^{N+1})}\|f\|_{\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)}[\omega]_{*}\Big{]}\omega(|x-y|).$
By the condition $\omega(\infty)=\infty$, we choose $N$ large enough and bound the bracketed term from above by $\leq 2\varepsilon[\omega]_{*}.$ Repeating the same proof with $y$ in place of $x$, we control the right sum on the line (33) and thus obtain $|f(x)-f(y)|\leq 4\varepsilon[\omega]_{*}\omega(|x-y|),$ that is, $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y).$ Then, we show that $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y)\subset\mathrm{VMO}^{\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y).$ Let $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y)$, $\varepsilon>0,$ and $R>0$ be such that $\operatorname{osc}^{\omega}_{\delta}(f)\leq\varepsilon,$ if $\delta\geq R.$ Let $M\gg R$ be a large constant, which we will specify later, and $Q\in\mathcal{Q}$ a cube such that $\ell(Q)\geq M.$ For every $y\in Q,$ denote by $Q_{R}(y)$ the cube centered at $y$ and of $\operatorname{diam}(Q_{R}(y))=R.$ Then, we have
@@ 3. Approximation in infinite dimensional spaces
In this section we study smooth approximations in Banach spaces with respect to the $\dot{C}^{0,\omega}$ seminorm. In Subsections 3.1 and 3.2 we establish results for real-valued functions $f:X\to\mathbb{R}$, where the approximating functions are not only $C^{k}$ or $C^{1,\alpha}$ smooth but also Lipschitz and with bounded support. In the last Subsection 3.3 we obtain $C^{\infty}$ smooth approximations of Banach-valued mappings $f:c_{0}(\mathcal{A})\to Y,$ that are not necessarily Lipschitz, but have bounded support and belong to $\dot{\operatorname{VC}}^{0,\omega}(X,Y).$ As a corollary, we deduce $C^{\infty}$ and $\dot{\operatorname{VC}}^{0,\omega}$ and approximations with bounded support for all $\dot{\operatorname{VC}}^{0,\omega}$ functions defined on $\mathbb{R}^{n}$ and with values in any Banach space.